Determine (a) the peak frequency deviation, (b) minimum bandwidth, and (c) baud for a binary FSK signal with a mark frequency of 38 kHz, a space frequency of 40 kHz, and an input bit rate of 4 kbps.


1
Expert's answer
2022-01-18T10:23:46-0500

a)

"2\Delta\>f=|fm-f_s|"

"\Delta\>f=\frac{(40-38)}{2}=1kHz"


b)

"B_{min}=2(\Delta\>f+f_b)"

"=2(1+4)"

"=10kHz"


c)

Baud for a binary FSK "=\frac{f_b}{N}" where "N=1"


"=\frac{4000}{1}=4000"